OpenStudy (anonymous):

Why is that when we change decimal to percent,we just move the decimal point to 2 places to the right then add a percent sign?

OpenStudy (skullpatrol):

Percent means "divided by 100"

OpenStudy (anonymous):

The simple idea is what skull said. So this explanation may be kind of messy, but maybe it explains it? Maybe it doesnt, but here goes :) You can think of a percent sign as meaning the same thing as \(\frac{1}{100}\). So if I give you 84%, it's like saying the same thing as \(84\cdot\frac{1}{100} = .84\). (dividing by 100 is what you're doing when moving 2 decimal places to the left). When going in the other direction, moving the decimal two places to the right, we're just going backwards. If I want to move the decimal place two spots to the right, I would normally have to multiply by 100. \(.84\cdot100\) But as you likely know, I cannot simply do this in math, I have to have equality. Add to both sides, times both sides, etc, there has to be equality in order for me to just perform an operation. Basically, I cannot multiply by 100 (I cannot move two decimal places to the right), unless I divide by 100 at the same time. So that means I have to do: \(.84\cdot\frac{100}{100}\) Now that I am allowed to do. The multiplication by 100 moves the decimal two places to the right, giving me: \(.84\cdot\frac{100}{100} =\frac{84}{100}\) But dividing by 100 is the same thing as having a percent sign, so I can then say: \(\frac{84}{100} = 84\%\)

OpenStudy (skullpatrol):

Any questions @Jaspercai ?